# A Survey on Unlearning in Large Language Models

### v1

@@ 5.2.1. Unlearning in Specialized Architectures and Scenarios
The field is moving towards addressing unlearning in sophisticated model architectures. Cheng and Amiri (2025) pioneer this effort for tool-augmented large language models (LLMs) by proposing ToolDelete, the first unlearning framework designed to remove a specific “skill” or the ability to use a particular tool, and they introduce a new membership inference attack (MIA) for evaluation. Similarly, the unique structure of Mixture-of-Experts (MoE) models presents a distinct challenge. Zhuang et al. (2025) find that unlearning a single expert is insufficient and propose the Selected Expert Unlearning Framework (SEUF) to effectively perform unlearning on MoE models. These works demonstrate that effective unlearning requires bespoke algorithms tailored to a model’s specific architecture and knowledge organization.
@@ 5.2.2. Unlearning as Tools
Unlearning is not only a goal in itself, but also a powerful tool when we further expand the scope of unlearning targets. Firstly, when choosing injected trojans or backdoor triggers as the target, unlearning can be an effective tool in defense (Jiang et al., 2025a; Zhao et al., 2025; Hernandez, 2024). Similarly, an opposite target, such as removing the safety alignment or disrupting the subsequent fine-tuning process on a base model, can convert unlearning to a means of attack (Rashid et al., 2025). Furthermore, if unlearning the selected training data and examining the changes in the model before and after unlearning, we can have novel insights into how different data components contribute to and influence the final model capabilities (Zhao et al., 2024; Isonuma and Titov, 2024). A powerful and accurate unlearning method will play an important role as a tool.
@@ 5.2.3. Unlearning beyond Data
Most existing studies focus on unlearning specific data instances. However, in practical scenarios, unlearning requests often target not only concrete data but also abstract concepts or capabilities, such as erroneous reasoning patterns, harmful ethical values, or unsafe skills (Lang et al., 2025; Li et al., 2024). Extending unlearning beyond the data to encompass abstract constructs is essential to prevent the propagation of incorrect or harmful knowledge. Achieving this goal may present two main pathways: one is to precisely identify and modify parameters or representations associated with particular concepts or abilities; the other leverages established alignment techniques, such as reinforcement learning, by designing appropriate reward mechanisms that penalize the generation of undesirable content.
@@ 5.2.4. Robust Unlearning
In light of the observed fragility of LLM unlearning, a significant research direction aims to develop techniques that enhance its robustness and long-term stability. These defensive efforts pursue two primary objectives: (1) to ensure that knowledge removal is thorough and persistent, thereby resisting attempts at recovery; and (2) to prevent the unlearning procedure from introducing new vulnerabilities or unintended side effects. Several existing studies address the first objective through robust unlearning frameworks (Tamirisa et al., 2025; Zhang et al., 2024a; Huu-Tien et al., 2025; Fan et al., 2025) or methods that strengthen the robustness of unlearned models (Zhang et al., 2024a; Huu-Tien et al., 2025). Nevertheless, given the proliferation of advanced attacks, achieving truly robust unlearning remains a critical and ongoing topic.
@@ 6. Conclusions
Acknowledgments
This work is partially supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (XDB0680201), the Beijing Natural Science Foundation (QY23156, 4252023).

### v2

@@ 5.2.1. Unlearning in Specialized Architectures and Scenarios
The field is moving towards addressing unlearning in sophisticated model architectures. cheng2025tool pioneer this effort for tool-augmented large language models (LLMs) by proposing ToolDelete, the first unlearning framework designed to remove a specific “skill” or the ability to use a particular tool, and they introduce a new membership inference attack (MIA) for evaluation. Similarly, the unique structure of Mixture-of-Experts (MoE) models presents a distinct challenge. zhuang2025seuf find that unlearning a single expert is insufficient and propose the Selected Expert Unlearning Framework (SEUF) to effectively perform unlearning on MoE models. These works demonstrate that effective unlearning requires bespoke algorithms tailored to a model’s specific architecture and knowledge organization.
@@ 5.2.2. Unlearning as Tools
Unlearning is not only a goal in itself, but also a powerful tool when we further expand the scope of unlearning targets. Firstly, when choosing injected trojans or backdoor triggers as the target, unlearning can be an effective tool in defense (jiang2025backdoor; zhao2025unlearning; hernandez2024if; Kazemi2024Unlearning; Lu2024Eraser). Similarly, an opposite target, such as removing the safety alignment or disrupting the subsequent fine-tuning process on a base model, can convert unlearning to a means of attack (Rashid2024Forget). Furthermore, if unlearning the selected training data and examining the changes in the model before and after unlearning, we can have novel insights into how different data components contribute to and influence the final model capabilities (zhao2024deciphering; isonuma2024unlearning). A powerful and accurate unlearning method will play an important role as a tool.
@@ 5.2.3. Unlearning beyond Data
Most existing studies focus on unlearning specific data instances. However, in practical scenarios, unlearning requests often target not only concrete data but also abstract concepts or capabilities, such as erroneous reasoning patterns, harmful ethical values, or unsafe skills (lang2025beyond; Li2024WMDP). Extending unlearning beyond the data to encompass abstract constructs is essential to prevent the propagation of incorrect or harmful knowledge. Achieving this goal may present two main pathways: one is to precisely identify and modify parameters or representations associated with particular concepts or abilities; the other leverages established alignment techniques, such as reinforcement learning, by designing appropriate reward mechanisms that penalize the generation of undesirable content.
@@ 5.2.4. Robust Unlearning
In light of the observed fragility of LLM unlearning, a significant research direction aims to develop techniques that enhance its robustness and long-term stability. These defensive efforts pursue two primary objectives: (1) to ensure that knowledge removal is thorough and persistent, thereby resisting attempts at recovery; and (2) to prevent the unlearning procedure from introducing new vulnerabilities or unintended side effects. Several existing studies address the first objective through robust unlearning frameworks (tamirisa2024tamper; zhang2024unforgettable; huu2025improving; fan2025towards) or methods that strengthen the robustness of unlearned models (zhang2024unforgettable; huu2025improving). Nevertheless, given the proliferation of advanced attacks, achieving truly robust unlearning remains a critical and ongoing topic.